OpenStudy (anonymous):

What is the value of the expression 11 + (fraction 1 over 2)4 ⋅ 48?

OpenStudy (alexandervonhumboldt2):

11+1/2*4*48=11+4/2*48=11+2*48=???

OpenStudy (anonymous):

?